OpenStudy (anonymous):

Completeing the square x^2+6x=7

OpenStudy (anonymous):

Dont know how to work it out at all. need help step by step please!

OpenStudy (anonymous):

u want to factor it right

OpenStudy (anonymous):

I dont know all i know is i am suppose to complete the square root.

OpenStudy (anonymous):

k w8

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

saifoo help him

OpenStudy (anonymous):

i don't know how to teach this one

OpenStudy (anonymous):

Blahh haah thanks me either. I wanst there for this so im trying to do it on my own no notes or anything !

OpenStudy (saifoo.khan):

simply use the quadratic formula.

OpenStudy (anonymous):

Thats all ?

OpenStudy (saifoo.khan):

Yes! =D

OpenStudy (saifoo.khan):

that is less confusing.

OpenStudy (anonymous):

Do you move the 7 over and make it -7?

OpenStudy (anonymous):

I got -6 +-Square root symbol 64/ 2

myininaya (myininaya):

\[x^2+6x=7\] they are asking you solve for x by completing the square \[x^2+6x+(\frac{6}{2})^2=7+(\frac{6}{2})^2\] \[(x+\frac{6}{2})^2=7+\frac{36}{4}\] \[(x+3)^2=\frac{28+36}{4}\] \[(x+3)^2=\frac{64}{4}\] now we are trying to get x by itself so we need to undo the square by square rooting both sides \[\sqrt{(x+3)^2}=\sqrt{\frac{64}{4}}\] but remember \[\sqrt{(x+3)^2}=|x+3|\] this gives two solutions \[x+3=\pm \sqrt{\frac{64}{4}}\] now subtract 3 on both sides \[x=-3 \pm \sqrt{\frac{64}{4}}=-3 \pm \frac{\sqrt{64}}{\sqrt{4}}=-3 \pm \frac{8}{2}=-3\pm 4\]

myininaya (myininaya):

-3+4=1 -3-4=-7

OpenStudy (anonymous):

Well never mind. Myininaya beat me too it. :)

OpenStudy (anonymous):

Thanka anyways. so uhmm what is that -3+4=1 and -3-4=_7?

OpenStudy (anonymous):

thanks myininaya i had for got the second step.........i had seen this after along time u brushed off all the rust from my mind........thanks a ton!!

myininaya (myininaya):

lol you're welcome i honestly don't understand what you mean what is -3+4=1 and -3-4=-7

OpenStudy (anonymous):

Where did that come from why did you put that down?

myininaya (myininaya):

i didn't have to i could have left the simplification to you

myininaya (myininaya):

so you don't understand how to get this \[-3\pm 4 \] or you don't know why -3+4=1 and why -3-4=-7?

OpenStudy (anonymous):

oh never mind i know what you did now i was confused on where you got the 3+4 and 3-4 but i see where you got it now

OpenStudy (anonymous):

you got it from the simplifyed answer sorry i had a moment

myininaya (myininaya):

lol you know \[-3 \pm 4 => -3+4 or -3-4\]

OpenStudy (anonymous):

Yeah ahah sorry i totally forgot that that is what you do.

myininaya (myininaya):

this is two expressions written as one

OpenStudy (anonymous):

Yeah gottcah. (: thanks you helped ALOTTT! Thanks soo much

myininaya (myininaya):

you could hav written it as ax^2+bx+c=0 and use the quadratic formula \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}=> x=\frac{-b -\sqrt{b^2-4ac}}{2a} \text{ or } x=\frac{-b + \sqrt{b^2-4ac}}{2a}\]

myininaya (myininaya):

but as i pointed out earlier it does say solve for completing the square i mainly just wanted to show you that the quadratic formula is an equation that provides two solutions

OpenStudy (anonymous):

Yeah thats whta i am doing and using the formula .